A triangular building with each side 30 m is 45 m tall. Find the cost of painting the building from outside at the rate of rupees 25 per square yard and the cost of fixing glass on the top of the roof at a cost of 200 rupees per square foot?

Answer by addingup(3677)

{[3(30*45)]/1.96}*25 = cost of painting the building on the outside. Use your calculator.
And the top is an equilateral triangle, each side is 30m:
A = (sqrt3/4)*a^2
A = [sqrt(3/4)]*30^2
A = 0.433*900 = 389.7 meters is the area of the triangular rooftop.
(389.7*10.764)*200 is the cost of fixing the glass on the roof. Use your calculator.